How many grams of H2O are needed to react with 3250 g CO? 


1
Expert's answer
2021-10-07T03:48:14-0400

The reaction is as shown below;

"H_2O+CO\to CO_2+H_2"

Molar mass of "CO=12+16=28"

Number of moles of CO="3250\/28=116.07moles"

Molar mass of "H_2O=2+16=18"

Mole ratio, "H_2O:CO=1:1"

Grams of "H_2O" reacted"=116.07\times 18=2089.26g"